OpenStudy (anonymous):

can you check my answer please..i'm not sure its correct :S..posted in replies

OpenStudy (anonymous):

\[\frac{ 3 }{ x+1 }=\frac{ 2 }{ x-1 }\] 2(x-1)=3(x+1) 2x-2=3x+3 2x-3x=2+3 -x=5 x=-5

OpenStudy (lgbasallote):

you multiplied wrong...

OpenStudy (lgbasallote):

should be 3(x-1) = 2(x+1)

OpenStudy (lgbasallote):

what you did was 3(x+1) = 2(x-1)

OpenStudy (anonymous):

so the answer would be -5 instead?

OpenStudy (lgbasallote):

3(x-1) = 2(x+1) 3x - 3 = 2x + 2 3x - 2x = 3 + 2 no i dont think it's -5

OpenStudy (anonymous):

so its still positive 5?

OpenStudy (lgbasallote):

yes it's +5 why "still"?

OpenStudy (anonymous):

dont worry i'm getting muddled..

OpenStudy (lgbasallote):

ah i see

OpenStudy (anonymous):

thankyou for your help..

OpenStudy (lgbasallote):

welcome